OpenStudy (3psilon):

If on of the roots of mx^2 + (4m-1)x + k = 0 is -1-2i, where m and k are real numbers. What is the value of K. I'm getting the equation for -1-2i but I keep getting k = 5 rather than the correct answer k= 5/2. PLease help! Thanks !

hartnn (hartnn):

(x+1)^2 -(2i)^2 x^2+2x+1 + 4 x^2+2x+5 so 4m-1 = 2 wait.

hartnn (hartnn):

u got x^2+2x+5=0 right ?

OpenStudy (3psilon):

yes

hartnn (hartnn):

there is another longer method, since -1-2i is the root, it must satisfy the equation so put x=-1-2i in mx^2 + (4m-1)x + k = 0 what u get ? then equate real and imaginary parts

OpenStudy (3psilon):

How do you equate real and imaginary parts?

hartnn (hartnn):

1st did u put x=-1-2i in mx^2 + (4m-1)x + k = 0

OpenStudy (3psilon):

Oh no I did not

OpenStudy (3psilon):

I plugged it in but now what do I do?

hartnn (hartnn):

mx^2 + (4m-1)x + k = 0 m(-1-2i)^2 + (4m-1)(-1-2i) + k = 0 m(1-4-4i) + (-4m-8mi+1+2i) + k=0 can u write next few steps of simplification ?

OpenStudy (3psilon):

-7m-8mi+1+2i +k

hartnn (hartnn):

-4mi -8mi = -12 mi

hartnn (hartnn):

so u get -7m-12mi+1+2i +k=0

OpenStudy (3psilon):

Ok ok then what do we do?

hartnn (hartnn):

(-7m+1+k) + i (-12m+2) = 0 real part = -7m+1+k = 0 imaginary part = -12m+2 = 0 <----find m from here

OpenStudy (3psilon):

m(-1-2i)^2 wouldn't it be m(-3+4i) ? you put m(1-4-4i)

hartnn (hartnn):

oh wait! my bad, so sorry. its m(-3+4i) only.

hartnn (hartnn):

-7m-4mi+1+2i +k=0 u get -4m+2=0 <----find m from here

hartnn (hartnn):

did u get m= 1/2 ? then put that in -7m+1+k = 0 to get k=5/2

hartnn (hartnn):

ask if u did not understand any step.

OpenStudy (3psilon):

I see what you did! Thank you hartnnn I'm getting the hang of this

hartnn (hartnn):

welcome ^_^ did u understand equating real and imaginary part ?

OpenStudy (3psilon):

Yes! Cause a+bi right

hartnn (hartnn):

yes, if a+bi =0 = 0+0i then a=0 and b=0

OpenStudy (3psilon):

Thanks man! Very very helpful :)

OpenStudy (3psilon):

You're awesome at math. Hope to be like you one day lol

hartnn (hartnn):

with practice , you can become better than me :)

OpenStudy (3psilon):

:) have a good night or day man

hartnn (hartnn):

same to u :)